# Search for Axion-Like Particles in High-Magnetic-Field Pulsars with NICER

### v1

@@ Abstract
Axion-like particles (ALPs) can couple to photons in strong magnetic fields, producing characteristic fluctuations in X-ray spectra. Using data from NASA’s Neutron Star Interior Composition EXplorer (NICER), We analyzed three pulsars, PSR J2229+6114, PSR J1849-0001, and PSR B0531+21, to search for such features. Each spectrum was modeled with a sliding-window power-law fitting method to identify local deviations from the smooth continuum. From these analyses, we derived constraints on the axion–photon coupling constant $g_{a\gamma\gamma}$ within a refined parameter space compared to previous studies, obtaining upper limits in the range $10^{-12}-10^{-14}GeV^{-1}$.
1 Introduction
The Standard Model (SM) of particle physics has achieved remarkable experimental success over the past several decades, yet several unresolved issues remain, among which the strong CP problem in quantum chromodynamics (QCD) is particularly compelling. This problem arises from the CP-violating term in the QCD Lagrangian:
$\mathcal{L}_{CP-vio}=\frac{\alpha_{s}}{4\pi}\theta\operatorname{Tr}\left(G_{\mu v}\tilde{G}^{\mu v}\right)$ (1)
where $G_{\mu\nu}$ is the gluon field strength tensor and $\tilde{G}^{\mu\nu}$ its dual. To solve this, Steven Weinberg [1] and Frank Wilczek [2] pproposed the axion: a pseudoscalar Nambu–Goldstone boson arising from the Peccei–Quinn mechanism [3]. Extensions of the SM predict more general axion-like particles (ALPs) that couple to photons via
@@ 1 Introduction
where $a$ is the axion field, $F$ and $\tilde{F}$ are the electromagnetic field-strength tensor and its dual, E and B are the electric field and magnetic field. This property enable photon-to-ALPs conversion in strong magnetic fields, the photon-ALP conversion probability in magnetic field is[4, 5].
$\mathcal{P}(\gamma\rightarrow a)=\mathcal{P}(a\rightarrow\gamma)=(\frac{{g_{a\gamma}}B_{e}}{q})^{2}{\sin^{2}(\frac{qL}{2})}$ (3)
where the $q=\frac{{m_{a}}^{2}-{\omega_{pl}}^{2}}{2E}$ is the momentum transfer, $B_{e}$ the effective magnetic field, and $L$ the photon path length that is the radius of pulsar here.
ALPs are probed by experiments targeting either non-relativistic or relativistic regimes. Non-relativistic ALPs, with low velocities are primary dark matter candidates, detected by haloscopes like Axion Dark Matter eXperiment (ADMX) [6, 7, 8, 9] in resonant cavity as well as constrained by cosmological evolution like Big Bang Nucleosynthesis (BBN) [10] and cosmic background [11] and so on.Relativistic ALPs are explored by helioscopes like CERN Axion Solar Telescope (CAST) [12], by light-shining-through-walls (LSW) experiments like Any Light Particle Search (ALPS)[13], and by pulsarscope studies [14].
@@
We assume the magnetic field distribution is uniform in each pulsar when we do the evaluation of probability. The magnetic field structure of magnetosphere is nonuniform certainly, likely dominated by a dipolar or multi-polar configuration. The uniform filed assumption thus serves as an effective model for estimating the characteristic conversion scale and the derived probabilities should be regarded as order-of-magnitude sensitivity rather than detailed prediction.
2 Data
We present the preliminary pulsar spectrum from NICER mission which is an external payload on the International Space Station (ISS) dedicated to high-precision soft X-ray timing and spectroscopy, installed in 2017. It’s scientific goals include enhancing the sensitivity of thermal and non-thermal emissions of neutron stars in the soft X-ray band, probing their interior composition, and investigating the origins of their dynamic phenomena. [18]
All data (see the Table 1 in Appendix) in this work were processed using HEASoft version 6.35.2 and NICER analysis software NICERDAS version 11. We applied calibration and event filtering with NICERL2, followed by spectral extraction with NICERL3-spect. This pipeline produces the necessary data products, including cleaned event spectra, background spectra, and instrument response files, which are then used for spectral analysis and plotting. For the background, we used ”3c50” [19] to be our background model.
The surface magnetic field of pulasrs is estimated under the standard magnetic dipole braking model [20], and the magnitudes are $2.03\times 10^{12}$ G for PSR J2229+6114 [21], $7.5\times 10^{11}$ G for PSR J1849-0001[22], and $1.2\times 10^{12}$ G for PSR B0531 (Crab) [23].For spectral analysis, we restricted range which the signal-to-noise ratio (SNR) should larger than 3.
3 Analysis
Each pulsar spectrum was modeled with an absorbed power-law function,
$F(E)=AE^{-\Gamma}$ (4)
to represent the underlying continuum emission. Figure 1 shows the spectra which we used and yellow area is the energy range for the analysis. To search for localized spectral deviations from the smooth baseline, we applied a sliding-window fitting approach. For energies below 2 keV, a window width of 4 bins and a step size of 2 bins were used; for energies above 2 keV, the window width and step size were set to 10 and 5 bins, respectively.
@@ 3 Analysis
Due to the overlapping nature of the windows, each energy bin is included in multiple fits and thus associated with more than one set of best-fit parameters. To determine the most representative local fit, we adopted the minimum $\chi^{2}$ criterion, selecting for each energy bin the fit corresponding to the lowest reduced $\chi^{2}$ value among all overlapping windows.
In the small-mixing limit($qL\ll 1$), the conversion probability can be rewrote to
$\mathcal{P}(\gamma\rightarrow a)\cong(\frac{g_{a\gamma B_{e}L}}{2})^{2}$ (5)
For a typical pulsar radius between 10 to 13 km, we take $L=13$ km when estimating the conversion scale.
To quantified deviations between data and fitted baseline, we compute the pull
$z_{i}=\frac{y_{data,i}-y_{model,i}}{\sigma_{i}s_{i}}\ \ \ \rm where\ s_{i}=\sqrt{\frac{\chi^{2}}{d.o.f}}$ (6)
and corresponding two-tailed Gaussian probability:
$p_{two-tailed,i}=2[1-\Phi(|z_{i}|)]$ (7)
where $\Phi$ is the Cumulative Distribution Function, which is used to represent random fluctuation would push the measurement at least this far away from prediction, This probability gives the chance that random noise would be produce a deviation ar least as large as $|z_{i}|$. To quantity the photon-Axion conversion probability at each energy bin, we define
$P_{conversion,i}=1-p_{two-tailed,i}$ (8)
In this formulation, $P_{conversion,i}$ gives the complementary probability that the observed deviation is not consistent with random fluctuations and can therefore be interpreted as the empirical conversion probability derived from data.
4 Result
We analyzed the NICER spectra of three pulsars, PSR J2229+6114, PSR J1849-0001 and PSR B0531+21. No statistically significant signal was detected in the NICER spectra of PSR J2229+6114, PSR J1849-0001, and PSR B0531+21. We therefore derive upper limits on the axion–photon coupling constant $g_{a\gamma\gamma}$. The most stringent constraint, $g_{a\gamma\gamma}=9.2\times 10^{-14}\,\mathrm{GeV^{-1}}$, arises from PSR J2229+6114, which has the highest surface magnetic field (B $\approx 2.0\times 10^{12}\,{\rm G}$). The results are consistent across alternative windowing schemes, indicating robustness against model-dependent effects.
Overall, both fitting schemes yield consistent parameter trends, conforming that the inferred residual features are insensitive to the specific window-selection criterion. Figure 2 shows the derived constraints for each pulsar and compares them with non-dark matter axion constraints and this result extends the existing bounds. Figure 2 shows the constraints compared with dark matter axion constraints from the astrophysical bounds.
@@ 4 Result
Figure 2: Axion-photon coupling constant constraints result from J2229+6114, J1849-0001 and B0531+21 pulsar spectra 2 for non-dark matter axion constraints and  2 for all dark matter constraints near X-ray range.
5 Conclusion
@@
We analyzed NICER observations of three rotation-powered pulsars to search for spectral signatures of photon–ALP conversion. Using a sliding-window power-law analysis, we derived constraints of $g_{a\gamma\gamma}\sim 10^{-12}–10^{-14}\,\mathrm{GeV^{-1}}$, with PSR J2229+6114 providing the strongest limit. These findings highlight the potential of X-ray observations of strongly magnetized pulsars as sensitive probes of axion-like particles in the relativistic regime.
Acknowledgment
@@
Appendix A Appendix
Pulsar Name B (G) energy range(keV) ObsID exposure time(s)

### v2

@@ Abstract
Axion-like particles (ALPs) can couple to photons in strong magnetic fields, producing characteristic fluctuations in X-ray spectra. Using data from NASA’s Neutron Star Interior Composition EXplorer (NICER), we analyzed three pulsars, PSR J2229+6114, PSR J1849-0001, and PSR B0531+21, to search for such features. Each spectrum was modeled with a sliding-window power-law fitting method to identify local deviations from the smooth continuum. From these analyses, we derived constraints on the axion–photon coupling constant $g_{a\gamma\gamma}$ from dipole and quadrupole magnetic field distribution assumption, obtaining upper limits in the range $10^{-10}-10^{-13}GeV^{-1}$ within the 0.8-10 keV mass region.
I Introduction
The Standard Model (SM) of particle physics has achieved remarkable experimental success over the past several decades, yet several unresolved issues remain, among which the strong CP problem in quantum chromodynamics (QCD) is particularly compelling[1, 2, 3]. This problem arises from the theoretically allowed but experimentally unobserved CP-violating term in the QCD Lagrangian:
$\mathcal{L}_{\text{CP-vio}}=\frac{\alpha_{s}}{4\pi}\theta\operatorname{Tr}\left(G_{\mu v}\tilde{G}^{\mu v}\right)$ (1)
where $G_{\mu\nu}$ is the gluon field strength tensor and $\tilde{G}^{\mu\nu}$ its dual. To resolve this, the Peccei-Quinn mechanism was proposed, giving rise to the axion—a pseudoscalar Nambu-Goldstone boson. Extensions of the SM predict more general axion-like particles (ALPs) that couple to photons via
@@ I Introduction
where $a$ is the axion field, $g_{a\gamma\gamma}$ is the coupling constant, $F$ and $\tilde{F}$ are the electromagnetic field-strength tensor and its dual, E and B are the electric field and magnetic field. For an electromagnetic wave, the interaction specifically requires the electric field component parallel to the external magnetic field ($E_{\parallel}B_{\perp}$). This property enables photon-to-ALPs conversion in strong magnetic fields; the resulting photon-ALP conversion probability in a magnetic field is determined by the mixing of the photon and ALP states as they propagate through the magnetosphere [4, 5].
ALPs are probed by experiments targeting either non-relativistic or relativistic regimes. Non-relativistic ALPs, with low velocities are primary dark matter candidates, detected by haloscopes like Axion Dark Matter eXperiment (ADMX) and Taiwan Axion Search Experiment with a Haloscope (TASEH) [6, 7, 8, 9, 10] in resonant cavity as well as constrained by cosmological evolution like Big Bang Nucleosynthesis (BBN) [11] and cosmic background [12] and so on.Relativistic ALPs are explored by helioscopes like CERN Axion Solar Telescope (CAST) [13], by light-shining-through-walls (LSW) experiments like Any Light Particle Search (ALPS)[14], and by pulsarscope studies [15].
@@
II Formula
The evolution of the coupled photon-ALP system as it traverses the pulsar magnetosphere can be described by a Schrödinger-like equation:
$i\frac{d}{dz}\Psi(z)=\mathcal{M}\Psi(z),\quad\Psi(z)=\begin{pmatrix}A_{\perp}(z)\\ a(z)\end{pmatrix}$ (3)
where $A_{\perp}(z)$ and $a(z)$ represent the photon and ALP field amplitudes with path length $z$, respectively. The mixing matrix $\mathcal{M}$ accounts for the interaction between the states:
$\mathcal{M}==\omega\mathbb{1}+\begin{pmatrix}\Delta_{\gamma}(z)&\Delta_{a\gamma}(z)\\ \Delta_{a\gamma}(z)&\Delta_{a}(z)\end{pmatrix}$ (4)
In this framework, $\omega$ is the oscillation frequency in vacuum, $\Delta_{\gamma}$ and $\Delta_{a}$ are the refractive indices for photons and ALPs, and $\Delta_{a\gamma}(z)=\frac{1}{2}g_{a\gamma\gamma}B_{\perp}(z)$ represents the interaction strength mediated by the transverse magnetic field.
To quantify the photon-to-ALP conversion, we consider the probability $\mathcal{P}(\gamma\rightarrow a)$ of a photon emitted from the stellar surface ($z=0$) converting into an ALP as it propagates to infinity. Following standard magnetospheric emission models, we assume the X-ray photons originate at or near the surface, where the dense stellar interior is opaque to X-ray propagation. The momentum transfer between the photon and ALP is given by $q=\frac{m_{a}^{2}-\omega_{pl}^{2}}{2E}$ , where the $\omega_{pl}$ is the plasma frequency. In the small-mixing limit($q\approx 0$), where the conversion efficiency is low, the probability can be expressed in an integral form:
$\mathcal{P}(\gamma\rightarrow a)=$ $\left|\int_{0}^{\infty}\Delta_{a\gamma}(z)e^{iqz}dz\right|^{2}$
$\hskip-14.22636pt=\left|\int_{0}^{\infty}\frac{1}{2}g_{a\gamma\gamma}B_{\perp}(z)e^{iqz}dz\right|^{2}$
$\hskip-14.22636pt\approx\left|\int_{0}^{\infty}\frac{1}{2}g_{a\gamma\gamma}B_{\perp}(z)\right|^{2}\text{if }q\approx 0\;.$ (5)
For a standard dipole distribution, the transverse magnetic field component is $B_{\perp}(r)=B(r)\sin\theta$, where $\theta$ is the angle between the magnetic axis and the line of sight. The field scales as $1/r^{3}$. By substituting $r=R_{\text{NS}}+z$, where $R_{\text{NS}}$ is the neutron star radius, we anchor the coordinate system to the emission site at the surface ($z=0$). Under this approximation, the integral simplifies to a clean analytical form:
$\hskip-14.22636pt\mathcal{P}(\gamma\rightarrow a)=\left|\frac{1}{2}g_{a\gamma\gamma}\sin\theta B_{0}R_{NS}^{3}\int_{0}^{\infty}\frac{1}{(R_{NS}+z)^{3}}dz\right|^{2}$
$\hskip 28.45274pt=\left|\frac{1}{4}g_{a\gamma\gamma}\sin\theta B_{0}R_{NS}\right|^{2}\;.$ (6)
Beyond the dipole assumption, we can expand this to a general multipole form. Since there is no free current outside the neutron star, the magnetic potential $\Phi_{M}$ satisfies Laplace’s equation, $\nabla^{2}\Phi_{M}=0$. In spherical coordinates, requiring the field to vanish at infinity implies a radial scaling of $B\propto 1/r^{l+2}$ for a multipole of order $l$:
$B_{\perp}(r,\theta)=B_{0}\left(\frac{R_{\text{NS}}}{r}\right)^{l+2}\sin\theta$ (7)
Using this generalized field in the conversion integral yields the final probability for any multipole order $l$:
$\mathcal{P}(\gamma\rightarrow a)=\left|\frac{1}{2(l+1)}g_{a\gamma\gamma}\sin\theta B_{0}R_{\text{NS}}\right|^{2}\;.$ (8)
This generalized result allows us to evaluate the sensitivity of our constraints to different magnetospheric models, with $l=1$ representing a standard dipole field and $l=2$ representing the quadrupole field.
III Analysis
We analyzed the soft X-ray spectra of three rotation-powered pulsars—PSR J2229+6114, PSR J1849-0001, and PSR B0531+21—using data from NASA’s NICER mission [19]. All spectra (see the Table 1 in Appendix) in this work were processed using the HEASoft version 6.35.2 and NICERDAS version 11 pipeline, with event filtering performed via NICERL2 and spectral extraction through NICERL3-spect. For our background subtraction, we utilized the ”3c50” empirical model[20]. All of the spectra show in Fig.  1, the yellow banding region means the fit range we used. Following the magnetospheric emission models for young pulsars [21, 22], we assume the X-ray photons in the 0.8–10 keV range originate from the polar cap or the near-surface vacuum gaps. Because these emission heights are negligible compared to the conversion path length, and given the star’s opacity to X-rays, we define the starting point of our integral at the neutron star surface.
To identify potential spectral fluctuations indicative of photon-ALP conversion, each pulsar’s continuum was modeled with an absorbed power-law function: $F(E)=AE^{-\Gamma}$. A sliding-window fitting approach was applied to detect localized deviations from this smooth baseline. We used a window width of 4 bins for energies below 2 keV and increased this to 10 bins for higher energies to account for decreasing photon statistics. Due to the overlapping nature of the windows, each energy bin is included in multiple fits and thus associated with more than one set of best-fit parameters. To determine the most representative local fit, we adopted the minimum $\chi^{2}$ criterion, selecting for each energy bin the fit corresponding to the lowest reduced $\chi^{2}$ value among all overlapping windows, all calculation are calculated at the 95% confidence level. The statistical significance of deviations between the data and our baseline model was quantified using the ”pull” ($z_{i}$):
$z_{i}=\frac{y_{\text{data,}i}-y_{\text{model,}i}}{\sigma_{i}s_{i}},\quad s_{i}=\sqrt{\frac{\chi^{2}}{d.o.f}}$ (9)
where $s_{i}$ is the scaling factor from the local fit. From this, we calculated a two-tailed Gaussian probability
$p_{\text{two-tailed,}i}=2[1-\Phi(|z_{i}|)]$ (10)
[…]
@@ III Analysis
IV Result
For a typical pulsar radius between 10 to 13 km, we take $R_{\text{NS}}=13$ km when estimating the conversion scale. To refine our constraints, we incorporated pulsar-specific geometry by assuming tilt angles ($\theta$) based on existing multi-wavelength research. We adopted $\theta\approx 60^{\circ}$ for PSR B0531+21 based on its torus-jet morphology, $\theta\approx 45^{\circ}$ for PSR J2229+6114 from pulse profile fitting, and a conservative $\theta\approx 10^{\circ}$ for the nearly aligned PSR J1849-0001.[23, 24, 25, 26]
With different angle assumption mentioned above, the coupling constant distribution show in Fig. 2, and the combined upper bound is also showed in Fig. 2. A central component of our analysis is the intense magnetic environment of these targets; the surface magnetic fields ($B_{0}$) were estimated using the standard magnetic dipole braking model[27], yielding values of $2.03\times 10^{12}$ G for PSR J2229+6114 , $7.5\times 10^{11}$ G for PSR J1849-0001, and $1.2\times 10^{12}$ G for PSR B0531+21.[28, 29, 30]
[…]
@@ IV Result
Figure 2: Axion-photon coupling constant constraints result from J2229+6114, J1849-0001 and B0531+21 pulsar spectra 2 for different tilt angle assumption and 2 upper bound of all of the pulsar in two different magnetic field distribution.
[…]
Figure 3: Constraints on the axion-photon coupling constant $g_{a\gamma\gamma}$ derived from NICER observations of PSR J2229+6114, PSR J1849-0001, and PSR B0531+21. 3 The derived limits compared with existing non-dark matter axion constraints in the X-ray regime. 3 The combined upper bounds across the full ALP mass range, compared against various astrophysical and cosmological dark matter constraints. All limits are calculated at the 95% confidence level.
Figure 3 shows the derived constraints for each pulsar and compares them with non-dark matter axion constraints near X-ray range and this result might provide new extension of the existing bounds. Figure 3 shows the constraints compared with dark matter axion constraints from the astrophysical bounds.
V Conclusion
@@
In summary, We analyzed NICER observations of three rotation-powered pulsars, PSR J2229+6114, PSR J1849-0001 and PSR B0531+21, to search for spectral signatures of photon–ALP conversion. Using a sliding-window power-law analysis, we derived constraints of $g_{a\gamma\gamma}\sim 10^{-10}$ $10^{-13}\,\mathrm{GeV^{-1}}$, by assuming the angle of pulsar tilt and the magnetic field distribution. The PSR J2229+6114 and B0531+21 providing the strongest limit based on the assumptions. Though the study gives us the rough story about the coupling constant, the further study toward these pulsar is still necessary. These findings highlight the potential of X-ray observations of strongly magnetized pulsars as sensitive probes of axion-like particles in the relativistic regime.
VI Acknowledgment
@@
Appendix A Summary of Observations
Table 1: Summary of the NICER pulsar observations and magnetospheric parameters. The surface magnetic fields ($B_{0}$) are estimated using the standard dipole braking model [27], and the observation IDs (ObsID) represent the cleaned data segments used for spectral fitting.
Pulsar Name Surface Magnetic Field (G) Energy Range (keV) ObsID Exposure Time (s)